# Bayesian network 3D event reconstruction in the Cygno optical TPC for dark matter direct detection

### v1

@@ Abstract
The CYGNO experiment is developing a high-resolution gaseous Time Projection Chamber with optical readout for directional dark matter searches. The detector uses a helium-tetrafluoromethane (He:CF4 60:40) gas mixture at atmospheric pressure and a triple Gas Electron Multiplier amplification stage, coupled with a scientific camera for high-resolution 2D imaging and fast photomultipliers for time-resolved scintillation light detection. This setup enables 3D event reconstruction: photomultipliers signals provide depth information, while the camera delivers high-precision transverse resolution. In this work, we present a Bayesian Network-based algorithm designed to reconstruct the events using only the photomultipliers signals, yielding a full 3D description of the particle trajectories. The algorithm models the light collection process probabilistically and estimates spatial and intensity parameters on the Gas Electron Multiplier plane, where light emission occurs. It is implemented within the Bayesian Analysis Toolkit and uses Markov Chain Monte Carlo sampling for posterior inference. Validation using data from the CYGNO LIME prototype shows accurate reconstruction of localized and extended tracks. Results demonstrate that the Bayesian approach enables robust 3D description and, when combined with camera data, further improves the precision of track reconstruction. This methodology represents a significant step forward in directional dark matter detection, enhancing the identification of nuclear recoil tracks with high spatial resolution.
@@ 1 Introduction
The CYGNO experiment Amaro:2022gub , part of the international CYGNUS proto-collaboration Vahsen:2020pzb , is developing a high-resolution gaseous Time Projection Chamber (TPC) with optical readout, optimized for directional dark matter detection Deaconu:2017vam . The TPC uses a helium-tetrafluoromethane (He:CF4) gas mixture in the 60:40 ratio and atmospheric pressure, which allows for efficient ionization and scintillation bib:Fraga:2003uu ; Antochi:2018otx ; Miernik:2007cwz . Charged particles interacting with the gas create ionization tracks. The resulting electrons drift under a uniform electric field toward a triple Gas Electron Multiplier (GEM) Sauli1997 stage, where they are amplified and generate secondary light emission. This light is recorded by two complementary detection systems: an Active Pixel Sensor of type scientific Complementary MOS (APS-sCMOS) , which captures a high-resolution 2D projection of the event on the GEM plane, and photomultipliers (PMTs), which collect time-resolved scintillation light, providing information on the particle’s path along the longitudinal (drift) coordinate.
@@
While the camera provides detailed spatial information in the plane parallel to the GEM stack ($XY$), it lacks direct depth sensitivity, making it inherently a 2D imaging system. In contrast, PMTs provide time-resolved signals of the light emitted, allowing the reconstruction of the development of the track along the direction orthogonal to the GEM plane ($\Delta Z$). Moreover, since the intensity of the light collected by the PMTs depends on the emission point on the GEM plane, it is possible to infer the transverse ($XY$) position of the source as well. Therefore, by analyzing the PMT signals, a full 3D event reconstruction can be achieved, independent of the camera image.
@@
To achieve this, we develop a reconstruction algorithm based on probabilistic graphical models, namely Bayesian Networks (BN) pearlbook ; jensenbook ; DAgostini:2003syq ; kollerbook , that infers the $(X,Y)$ position of ionization tracks on the GEM plane from PMT signals, and estimates the light emitted during the amplification process, and thus reconstructing the particle’s energy. This information is combined with the $\Delta Z$ component extracted from the analysis of PMT waveforms, particularly their time profile. Once the 3D reconstruction from PMTs is obtained, it can be matched with the camera image, which provides superior $(X,Y)$ spatial resolution. This combination enables precise 3D reconstruction of the ionizing event. Merging these two independent measurements improves both spatial and energy resolution, significantly enhancing event characterization. This methodology marks an important advancement for directional dark matter detection, enabling precise identification of nuclear recoil tracks with improved spatial resolution.
@@ 2.1 The LIME prototype
The Long Imaging ModulE (LIME) LIME , shown in Fig. 1, is a prototype for the future CYGNO experiment. It consists of a box-shaped TPC with a drift distance of 50 cm and a transverse area of 33 $\times$ 33  cm2. The He:CF4 gas mixture is enclosed in a 10 mm-thick PMMA box, surrounded by a field cage composed of 34 copper rings with a cross-section of 330 mm $\times$ 330 mm, each 10 mm thick and spaced 4 mm apart. The field cage is bounded by a 0.5 mm-thick copper cathode on one side and a triple-GEM stack on the other. Each GEM has holes of 50 $\mu$m with a pitch of 140 $\mu$m and is separated by a 2 mm gap. A conical black PMMA structure is mounted on the side of the GEM stack to house a Hamamatsu ORCA-Fusion APS-sCMOS camera and four Hamamatsu R7378A PMTs, each with a 25.4 mm diameter and a quantum efficiency of about 25% in the 300–500 nm wavelength range. The CF4 emits light in two broad continua, peaked around 290 nm and 620 nm  FRAGA200388_rydberg_state , with the PMT sensitivity matching the UV component.The camera is equipped with a Schneider Xenon 0.95/25-0037 lens, featuring a focal length of 25.6 mm and an aperture ratio (f-number) of 0.95. It is positioned centrally in front of the GEM plane at a distance of 62.3 cm, while the four PMTs are located at the corners of a square plane parallel to the GEMs, 19 cm away from them. In this configuration, the camera’s field of view (FOV) covers an area of 35.7 cm $\times$ 35.7 cm, corresponding to a pixel granularity of 155 $\mu$m. Side and front views of the geometric arrangement between the camera, PMTs, and the GEM plane are shown in Figure 2.
@@
When a particle interacts within the LIME detector, it ionizes the gas mixture, and the resulting electrons drift toward the anode with a mean velocity of $v$ = 5.5 cm/$\mu$s under an electric field of 0.8 kV/cm. Upon reaching the GEM stage, the electrons undergo avalanche multiplication within the GEM holes, where the local electric field reaches up to 40 kV/cm  ELY . During this process, scintillation light is emitted primarily by the CF4 component of the gas. Part of the resulting photons are detected by the PMTs, enabling time-based reconstruction of the particle’s trajectory orthogonal to the GEM plane, and by the camera, which records the profile to the GEM plane.
Figure 3: Example of an event recorded with the LIME’s optical readout, illustrating (3(a)) the image acquired by the APS-sCMOS camera during a 300 ms exposure with four distinct tracks: three localized clusters and one extended ionization trail; and (3(b)) the PMT signals (inverted for clarity) recorded within the same acquisition window, each associated to one of the ionization in the picture.
An event in LIME consists of a combination of a camera image and a set of PMT signals. The data acquisition system (DAQ) operates as follows: the camera runs in triggerless mode with a 300 ms exposure, while PMT signals are continuously monitored. These signals pass through a leading-edge discriminator, and a trigger is issued if at least two PMTs exceed a predefined voltage threshold within the same time window. When a trigger occurs, the image is saved along with the corresponding PMT signals recorded over an 1.40 $\mu$s time window. In cases where multiple PMT triggers occur within the same camera exposure, several sets of PMT signals are associated with a single image. Figure 3 shows an example of such an event recorded in LIME, where multiple tracks are visible in the image – three short, localized tracks and one extended track – along with the corresponding PMT signals generated by the ionization processes. In these situations, the Bayesian inference algorithm can be employed to associate each set of waveforms with its corresponding track observed in the image. To analyze such events, the CYGNO collaboration developed a reconstruction algorithm Baracchini_2021 that identifies light clusters in the APS images and reconstructs their physical properties, including shape, light intensity, and direction on the GEM plane. The analysis of PMT signals has been developed in parallel with the present work and will be discussed in detail in a forthcoming publication.
@@ 2.2 PMTs signals
The signal in the $i$-th PMT corresponds to a voltage $V_{i}(t)$, as shown in Fig. 3(b). The total charge $Q_{i}$ collected by the PMT is proportional to the incident light $L_{i}$, and can be expressed as:
@@
where $\xi_{i}$ is a calibration factor (which may vary for each PMT), $\mathcal{R}=50\,\rm{\Omega}$ is the termination impedance and $\Delta t$ is the integration interval that will be specified later in the paper.
The relationship between the total light $L_{j}$ emitted in the position $(X_{j},Y_{j})$ of the GEMs and the amount $L_{ij}$ collected by the $i$-th PMT is modeled assuming Lambertian emission, i.e., radiation from a perfectly diffuse source optics . The Lambertian emission model provides a simplified yet effective description of the GEM light emission. In this scenario, the radiance of the emitting surface (i.e., the power emitted per unit solid angle per unit projected area) is independent of the viewing angle. The radiant power received at a distance $R_{ij}$, as illustrated in Fig. 4, is then given by:
$\Phi=\frac{L_{j}A_{j}A_{i}\cos^{2}{\theta_{ij}}}{R_{ij}^{2}}$ (2)
where $A_{j}$ and $A_{i}$ are the areas of the light source and the PMT, respectively. Noting that $\cos{\theta_{ij}}=h/R_{ij}$, where $h$ is the distance between the PMT and the GEM planes, the expression simplifies to:
$L_{ij}\propto\frac{L_{j}}{R_{ij}^{4}}$ (3)
This equation is the well-known cosine-fourth-power law of illumination. Given that this is an approximation optics , and thus might not fully capture all subtle details of the actual light emission characteristics, dedicated measurements validated this approximation: by fitting the parameter $\alpha$ in a generalized emission model $L_{ij}\propto L_{j}/R_{ij}^{\alpha}$, we obtained $\alpha=4.0$, with a total uncertainty within 10%. This accuracy is adequate for the purposes of our reconstruction algorithm and practical applications. Finally, the photocathode angular response of the PMTs is neglected, as it is minimal at the viewing angles relevant to the experiment pmt_handbook .
3 Application of Bayesian Networks for PMT-Based 3D Reconstruction
In Bayesian inference, Bayes’ theorem is used to update the probability of a model or a set of parameters as new evidence or information becomes available, in the form of experimental observations. The joint posterior probability $p(\bm{\theta}|\{x\})$ derives from the likelihood $p(\{x\}|\bm{\theta})$ and the prior probability $\pi(\bm{\theta})$ as:
$p(\bm{\theta}|\{x\})=\frac{p(\{x\}|\bm{\theta})\cdot\pi(\bm{\theta})}{p(\{x\})},$ (4)
where $\{x\}$ represents the data, and $\bm{\theta}$ denotes the parameters describing the experimental conditions or the theoretical assumptions.
In the case of the PMT-based reconstruction in the LIME prototype, the likelihood used in the inference process of Eq. 4 is defined as:
$\begin{split}p(\{x\}|\bm{\theta})&=\prod_{j=1}^{N}\prod_{i=1}^{4}\mathcal{N}(Q_{ij}\mid L^{\prime}_{ij}(\bm{\theta}))\\ &=\prod_{j=1}^{N}\prod_{i=1}^{4}\frac{1}{\sqrt{2\pi}\sigma_{ij}}\exp\left[-\frac{(Q_{ij}-L^{\prime}_{ij}(\bm{\theta}))^{2}}{2\sigma_{ij}^{2}}\right]\,,\end{split}$ (5)
where $j$ runs over all light-emitting sources, $i$ is the PMT index, and $Q_{ij}$ is the charge measured by the $i$-th PMT for the $j$-th emission point. The predicted signal $L^{\prime}_{ij}$ depends on the emitted light $L_{j}$, the emission position $(X_{j},Y_{j})$, the PMT position $(X_{i},Y_{i},h)$, and the PMT response, as defined by Eqs. 1 and 3:
$L^{\prime}_{ij}=C_{i}\,\frac{L_{j}}{R_{ij}^{\alpha}}\,,$ (6)
where $R_{ij}=\sqrt{(X_{j}-X_{i})^{2}+(Y_{j}-Y_{i})^{2}+h^{2}}$ is the distance between the source and the PMT, as shown in Fig. 4. The parameter $\alpha$ is fixed to 4, assuming a Lambertian emission profile. The observed charges $Q_{ij}$ are assumed to be independently normally distributed around the predicted values $L^{\prime}_{ij}(\bm{\theta})$, with standard deviation $\sigma_{ij}$ taken as 10% of $L^{\prime}_{ij}$. This relation allows us to define the physical dimensions of the calibration constants as $[C_{i}]=[L_{j}]$[L4][Q]. Here, $[L_{j}]$ denotes the physical dimension of the emitted light (e.g. number of photons), and [Q] corresponds to the physical dimension of the electric charge.
@@ 3 Application of Bayesian Networks for PMT-Based 3D Reconstruction
In Fig. 5 the graphical representation of the likelihood through a Bayesian network. The Bayesian network formalism is particularly useful as it makes the probabilistic assumptions explicit and allows flexible and modular extension of the model. This approach differs significantly from classical reconstruction strategies commonly used in optical TPCs, such as centroid estimators, clusterization algorithms, or $\chi^{2}$-based fitting methods. While those techniques are computationally efficient, they often lack a consistent way to propagate uncertainties and may struggle with complex detector geometries or signal topologies. In contrast, the Bayesian network encodes the underlying physics and geometry explicitly, enabling the direct computation of posterior distributions for all parameters of interest. Despite the use of sampling methods, the relatively small number of parameters per event makes the inference tractable. Moreover, the modular structure of the Bayesian model allows for transparent extensions to more complex scenarios such as multiple tracks, energy-dependent emission models, or prior-informed inference in low-signal regimes.
Figure 5: Bayesian network adopted in the PMT reconstruction. The node $Q_{ij}$ is the collected charge by $i$-th PMT for the $j$-th signal; $L^{\prime}_{ij}$ is the light reaching the PMT; $L_{j}$ is the light produced at the GEM hole; $X_{j}$ and $Y_{j}$ are the position of the emitted light in the GEM plane; $C_{i}$ are the proportionality factors relating light to charge; $X_{i}$ and $Y_{i}$ are the positions of the $i$-th PMT; $h$ is the distance between the GEM plane and the PMT plane; and $\alpha$ is the power dependence of the light reaching the PMTs on the distance. A solid arrow between nodes represents a probabilistic link between the two variables, while a dashed arrow between nodes represent a deterministic link between the two.

### v2

@@ Abstract
The CYGNO experiment is developing a high-resolution gaseous Time Projection Chamber with optical readout for directional dark matter searches. The detector uses a helium-tetrafluoromethane (He:CF4 60:40) gas mixture at atmospheric pressure and a triple Gas Electron Multiplier amplification stage, coupled with a scientific camera for high-resolution 2D imaging and fast photomultipliers for time-resolved scintillation light detection. This setup enables 3D event reconstruction: photomultiplier signals provide depth information, while the camera delivers high-precision transverse resolution. In this work, we present a Bayesian Network-based algorithm designed to reconstruct the events using only the photomultiplier signals, inferring a 3D description of the particle trajectories. The algorithm models the light collection process probabilistically and estimates spatial and intensity parameters on the Gas Electron Multiplier plane, where light emission occurs. It is implemented within the Bayesian Analysis Toolkit and uses Markov Chain Monte Carlo sampling for posterior inference. Validation using data from the CYGNO LIME prototype shows accurate reconstruction of localized and extended straight tracks. Results demonstrate that the Bayesian approach enables robust 3D description and, when combined with camera data, opens the way to future improvements in spatial and energy resolution. This methodology represents a significant step forward in directional dark matter detection, enhancing the identification of nuclear recoil tracks with high spatial resolution.
@@ 1 Introduction
The CYGNO experiment Amaro:2022gub, part of the international CYGNUS proto-collaboration Vahsen:2020pzb, is developing a high-resolution gaseous Time Projection Chamber (TPC) with optical readout, optimized for directional dark matter detection Deaconu:2017vam. The TPC uses a helium-tetrafluoromethane (He:CF4) gas mixture in the 60:40 ratio at atmospheric pressure, which allows for efficient ionization and scintillation bib:Fraga:2003uu; Antochi:2018otx; Miernik:2007cwz. Charged particles interacting with the gas create ionization tracks. The resulting electrons drift under a uniform electric field toward a triple Gas Electron Multiplier (GEM) Sauli1997 stage, where they are amplified and generate secondary light emission. This light is recorded by two complementary detection systems: an Active Pixel Sensor of type scientific Complementary MOS (APS-sCMOS), which captures a high-resolution 2D projection of the event on the GEM plane, and photomultipliers (PMTs), which collect time-resolved scintillation light, providing information on the particle’s path along the longitudinal (drift) coordinate. Although Timepix-based cameras could in principle provide nanosecond time resolution and hence a 3D reconstruction, they offer a much smaller number of pixels, higher per-pixel noise, and significantly higher costs. For large-area optical readouts, the APS-sCMOS combined with PMTs, as adopted in this work, remains more suitable.
@@
While the camera provides detailed spatial information in the plane parallel to the GEM stack ($XY$), it lacks direct depth sensitivity, making it inherently a 2D imaging system. In contrast, PMTs provide time-resolved signals of the light emitted, allowing the reconstruction of the development of the track along the direction orthogonal to the GEM plane ($\Delta Z$). Moreover, since the intensity of the light collected by the PMTs depends on the emission point on the GEM plane, it is possible to infer the transverse ($XY$) position of the source as well. Therefore, by analyzing the PMT signals, a 3D event reconstruction can be achieved, independent of the camera image.
In this work we focus on the 3D reconstruction of short straight tracks, which are the expected signature of low-energy nuclear recoils induced by dark matter interactions. Our present goal is to demonstrate robust reconstruction of this topology, which provides the basis for future extraction of the track angle and, ultimately, the direction of the incoming particle. Such directionality would offer a powerful handle to discriminate dark matter from background. For this reason, the reconstruction algorithm is optimized for straight-track topologies. More sophisticated approaches could in principle be developed to recover more complex event shapes, such as those induced by longer electronic recoils, but this lies beyond the scope of the present work and represents a promising direction for future studies.
Several collaborations pursue the development of gaseous TPCs for directional detection, such as NEWAGE NEWAGE2015, MIMAC MIMAC2011, DMTPC DMTPC2011, DRIFT DRIFT2012, all aiming to exploit track reconstruction to achieve sensitivity to the incoming WIMP direction. In addition to gaseous TPCs, R&D efforts have also explored the possibility of exploiting columnar recombination in liquid argon to gain directional sensitivity, as pursued in the RED program associated with DarkSide DarkSide20k:2025tes; Bondar2017. Directional TPCs have also been proposed for neutrino physics Lisotti:2024fco, and for neutrinoless double-beta decay NEXT:2012rto.
@@
To achieve a PMT-only 3D reconstruction, we develop a reconstruction algorithm based on probabilistic graphical models, namely Bayesian Networks (BN) pearlbook; jensenbook; DAgostini:2003syq; kollerbook, that infers the $(X,Y)$ position of ionization tracks on the GEM plane from PMT signals, and estimates the light emitted during the amplification process, thus reconstructing the particle’s energy. This information is combined with the $\Delta Z$ component extracted from the analysis of PMT waveforms, particularly their time profile. Once the 3D reconstruction from PMTs is obtained, it can be matched with the camera image, which provides superior $(X,Y)$ spatial resolution and an independent energy measurement. This combination enables precise 3D reconstruction of the ionizing event and represents a promising path to further enhance the energy resolution, which will be quantitatively demonstrated in a forthcoming paper. This methodology marks an important advancement for directional dark matter detection, enabling precise identification of nuclear recoil tracks with improved spatial resolution.
@@ 2.1 The LIME prototype
[…]
@@
[…]
Figure 3: Example of an event recorded with the LIME’s optical readout, illustrating (3(a)) the image acquired by the APS-sCMOS camera during a 300 ms exposure with four distinct tracks: two localized clusters; one extended straight ionization trail; and a curly scattered track (electron recoil). Figure (3(b)) shows the PMT signals (inverted for clarity) recorded within the same acquisition window, each associated to one of the ionization in the picture.
[…]
@@ 2.2 PMTs signals
The relationship between the light per unit surface per steradian $L_{j}$ emitted from position $(X_{j},Y_{j})$ of the GEMs and the corresponding light $L_{ij}$ collected by the $i$-th PMT is modeled under the assumption of Lambertian emission, i.e. radiation from a perfectly diffuse source with angle-independent radiance optics. In this case, the light received at a distance $R_{ij}$, as shown in Fig. 4, is
$L_{ij}=\frac{L_{j}A_{j}A_{i}\cos^{2}{\theta_{ij}}}{R_{ij}^{2}},$ (1)
where $A_{j}$ and $A_{i}$ denote the source and PMT areas. With $\cos\theta_{ij}=h/R_{ij}$, where $h=19$ cm is the fixed PMT–GEM separation, the expression reduces to
$L_{ij}\propto\frac{L_{j}}{R_{ij}^{4}},$ (2)
the cosine-fourth-power law of illumination. The PMT angular response was neglected, as it is modest at the viewing angles of interest pmt_handbook. The signal in the $i$-th PMT is a voltage $V_{ij}(t)$ (Fig. 3(b)). The corresponding charge $Q_{ij}$, proportional to $L_{ij}$, is
$Q_{ij}=\frac{1}{\mathcal{R}}\int_{\Delta t}V_{ij}(t)\,dt\,\propto L_{ij}$ (3)
where $\mathcal{R}=50\,\mathrm{\Omega}$ the termination impedance, and $\Delta t$ the integration window. Combining Eqs. 2 and 3 yields
$Q_{ij}=C_{i}\frac{L_{j}}{R_{ij}^{\alpha}},$ (4)
with $\alpha=4$ and $C_{i}$ calibration constants which include the relative gain and efficiency of $i$-th PMT as well as a global calibration factor (absolute scale) and the proportionality factors omitted in Eqs. 2, 3. Their dimensions are $C_{i}]=[\mathrm{L}]^{4}[\mathrm{Q}]/[L_{j}]$, with $L_{j}]$, $\mathrm{Q}]$, and $\mathrm{L}]$ the dimensions of radiance, charge, and length, respectively.
The PMT high voltages have been pre-adjusted with LED bench tests in order to equalize the individual gains. Subsequently, dataset of nearly mono-energetic, localized events is used to characterize the PMT response. Such events are obtained by exposing the detector to a 55Fe radioactive source, which emits $K_{\alpha}$ and $K_{\beta}$ X-rays from 55Mn. Due to the limited energy resolution, these lines merge into an effectively monochromatic peak at 5.9 keV. A sample of 55Fe events is selected following the procedure described in Refs. LIME; bib:fe55. The PMT waveforms produced by these electronic recoils typically show a single peak (Fig. 3). The integrated charge is computed within a 60 ns window centered on the peak. Events are further required to exhibit a single light spot on the GEM plane, as reconstructed by the optical camera, with well-defined positions $(X_{j},Y_{j})$.
Using this dataset and the $(X_{j},Y_{j})$ positions as reconstructed by the camera, we validate the cosine-fourth-power law of illumination by fitting the exponent $\alpha$ in the generalized form $L_{ij}\propto L_{j}/R_{ij}^{\alpha}$, obtaining $\alpha=4.0$, consistent with the model within less than 10%.
The PMT response is then characterized by comparing the measured charge $Q_{ij}$ with its expected value $\mu_{ij}$ defined in Eq. 4. Figure 5 shows, for each PMT, the relative standard deviation $\sigma_{ij}/\mu_{ij}$ as a function of $\mu_{ij}$. The dynamic range is sampled by interactions of 5.9 keV X-rays at different positions on the GEM plane. The insets display the charge dispersion for three selected points; in all cases, the distributions are well described by a Gaussian model.
Figure 5: Relative standard deviation, for each PMT, of the measured charge as a function of the expected value $\mu_{ij}$ from Eq. 4. The expected value spans over the dynamic range because of a 55Fe mono-energetic source occurring at different positions on the GEM plane. The colored points are the measured data, the black curve is the fit of the model described in Eq. 5 for all the PMTs and the blue band represents the $3\sigma$ uncertainty range of the inference process. The insets display the relative charge dispersion for the three points indicated by the arrows with the relative Gaussian model superimposed.
The general model adopted to describe the charge resolution is:
$\frac{\sigma_{ij}}{\mu_{ij}}=\frac{A}{\sqrt{\mu_{ij}}}+B$ (5)
where $A$ is the stochastic and $B$ the constant term accounting for systematic effects in the response model. In principle, a pair $(A_{i},B_{i})$ would be required to describe each PMT independently. However, given the comparable energy resolution observed across the photo-sensors, a single response function was found to adequately describe all of them. The fitted parameter values are:
$A=$ $(0.021\pm 0.002)\,\sqrt{\text{nC}}$
$B=$ $(0.06\pm 0.01)$
$\rho_{AB}=$ $-97\%$ (6)
We also tested an extended three-parameter fit including an additional term proportional to $\mu^{-1}_{ij}$} to account for charge-independent effects such as PMT electronic noise, which dominate at very low signals. However, no appreciable improvement is observed in the residuals, which remain equally well described by the two-parameter model. Since our study is limited to the same charge range shown in Fig. 5, the two-parameter description is both simple and effective, though it should not be regarded as a general measurement of the detector resolution outside this range.
3 Application of Bayesian Networks for PMT-Based 3D Reconstruction
In Bayesian inference, Bayes’ theorem is used to update the probability of a model or a set of parameters as new evidence or information becomes available, in the form of experimental observations. Denoting with $\{x\}$ the data, and with $\bm{\theta}$ the parameters describing the experimental conditions or the theoretical assumptions, the joint posterior probability $p(\bm{\theta}|\{x\})$ derives from the product of the direct probability $p(\{x\}|\bm{\theta})$, the prior probability $\pi(\bm{\theta})$ and the data probability $p(\{x\})$ as:
$p(\bm{\theta}|\{x\})=\frac{p(\{x\}|\bm{\theta})\cdot\pi(\bm{\theta})}{p(\{x\})}.$ (7)
When the direct probability $p(\{x\}|\bm{\theta})$ is regarded as a function of the parameters $\bm{\theta}$ for some given data $\{x\}$ it is referred to as the likelihood function $\mathcal{L}(\bm{\theta};\{x\})$.
In the case of the PMT-based reconstruction in the LIME prototype, the likelihood used in the inference process of Eq.  7 is defined as:
@@ 3 Application of Bayesian Networks for PMT-Based 3D Reconstruction
where $j$ runs over all light-emitting sources, $i$ denotes the PMT index, and $Q_{ij}$ is the charge measured by the $i$-th PMT for the $j$-th emission point. All parameters describing the model and the experiment are collected in $\bm{\theta}$: the geometry $(X_{i},Y_{i},h)$, which is considered known and kept fixed; the global exponent $\alpha$, also fixed; the calibration parameters $(C_{i})$, with priors taken from dedicated measurements; and the source variables $(X_{j},Y_{j},L_{j})$. The observed charges $Q_{ij}$ are assumed to be independently normally distributed around the expected values $\mu_{ij}(\bm{\theta})$ from Eq. 4, with standard deviation $\sigma_{ij}(\bm{\theta},A,B)$ described by Eq. 5. The parameters $A$ and $B$ of the resolution function are passed to the fit as nuisance parameters with the bivariate normal prior specified in Eq. 6.
Fig. 6 shows the graphical representation of the likelihood through a Bayesian network. Plates indicate repetition over events ($j$) and over PMTs ($i$); solid arrows indicate probabilistic links, dashed arrows deterministic ones; primordial nodes are either fixed (grey) or assigned a prior (white). Inference in this graphical model proceeds in both directions: once a node is observed, the corresponding information propagates backward through the network. For instance, measuring the charges $Q_{ij}$ allows us to infer the underlying event variables $X_{j}$, $Y_{j}$, and $L_{j}$, so that measurements constrain not only the directly connected nodes but also their ancestors.
@@
The Bayesian network formalism is particularly useful as it makes the probabilistic assumptions explicit and allows flexible and modular extension of the model. This approach differs significantly from classical reconstruction strategies commonly used in optical TPCs, such as centroid estimators, clusterization algorithms, or $\chi^{2}$-based fitting methods. While those techniques are computationally efficient, they often lack a consistent way to propagate uncertainties and may struggle with complex detector geometries or signal topologies. In contrast, the Bayesian network encodes the underlying physics and geometry explicitly, enabling the direct computation of posterior distributions for all parameters of interest. Despite the use of sampling methods, the relatively small number of parameters per event makes the inference tractable. Moreover, the modular structure of the Bayesian model allows for transparent extensions to more complex scenarios such as multiple tracks, energy-dependent emission models, or prior-informed inference in low-signal regimes.
Possible extensions in our model may also include additional uncertain parameters, such as the detector geometry $(X_{i},Y_{i},h)$ or the global exponent $\alpha$, which are currently treated as fixed: the geometry is precisely determined and its uncertainty is negligible, while for the exponent $\alpha$ we rely on the strong theoretical prior of the Lambertian model ($\alpha=4$), consistent with our data within $\sim$10%.
Figure 6: Bayesian network adopted in the PMT reconstruction. Each event $j\in N$ produces a light yield $L_{j}$ at position $(X_{j},Y_{j})$, while each PMT $i\in N_{PMTs}$ is characterized by a calibration coefficient $C_{i}$ and position $(X_{i},Y_{i})$. The expected mean signal $\mu_{ij}$ depends on these quantities, on global parameters $(\alpha,h)$, while $\sigma_{ij}$ depends also on nuisance terms $(A,B)$. The observed charge $Q_{ij}$ is modeled as a Gaussian distribution $\mathcal{N}$ with mean $\mu_{ij}$ and width $\sigma_{ij}$. Plates indicate repetition over events ($j$) and over PMTs ($i$); solid arrows indicate probabilistic links, dashed arrows deterministic ones. Primordial nodes are either fixed (grey) or assigned a prior (white).